Jhon is 6 years olden than his beother .he will be twice as old as his brother in 4 years.how old is jhon now?
Found 2 solutions by macston, FrankM:
Answer by macston(5194)

.
B=brother's age; J=john's age=B+6
.
J+4=2(B+4)
(B+6)+4=2(B+4)
B+10=2B+8
2=B
.
Brother is 2 years old.
.
J=B+6=2+6=8
.
ANSWER: John is 8 years old.
.
CHECK:
J+4=2(B+4)
8+4=2(2+4)
12=2(6)
12=12
.

Answer by FrankM(1040)

J=B+6
J+4=2(B+4)
B+6+4=2B+8
B+10=2B+8
10-8=2B-B
B=2, J=8 (Now)
In 4 years, B=6 J=12